OpenStudy (anonymous):

find the rectangular form of the equation r=2cosø A. x^2+y^2=2y B.2c^2+y^2=x C. X^2+Y^2=2 D none of the above

OpenStudy (jdoe0001):

tried by multiplying both sides by "r" yet?

OpenStudy (anonymous):

what do you mean ? like i want to take r=2cosø and multiply it by r

OpenStudy (jdoe0001):

\(\bf r=2cos(\theta)\implies {\color{brown}{ r^2}}=2{\color{brown}{ rcos(\theta)}}\qquad \begin{array}{llll} r^2\to x^2+y^2\\ \bf rcos(\theta)\to x \end{array}\)